# Disruptions of stars and binary systems on chaotic orbits in the axisymmetric Milky Way center

### v1

@@ 4.2.1 Lyapunov exponents
$\lambda$ should only be calculated for the section of a trajectory for which the displacement is growing exponentially. It is common to use the convention of $\lambda=0$ for trajectories which show no such exponential growth and thus are assumed to be regular. Following the scheme outlined in Vasiliev (2013), we can efficiently estimates $\lambda$ in units of the circular period $T_{\rm c}$, such that its value is normally between 0 and a maximum close to 1 (corresponding to an trajectory that diverges over a single orbit).
A completely general and robust quantification of $\lambda$ is a difficult task and we note two edge cases that infrequently occur. The first comes from demanding too high an integration accuracy from the integrator, which can result in ’ghost’ timesteps (with ${\rm d}t_{i}=t_{i+1}-t_{i}=0$) causing $\lambda$ to be undefined. The second comes from lower integration accuracy, leading to larger timesteps which are still well optimized for resolving the orbital path but which may now be insufficiently spaced for calculating $\lambda$, and thus give erroneous $\lambda>0$ for regular orbits (as identified, for example from the SoS).
@@ 4.3 Uniform filling of the chaotic region
A useful behaviour of low-$l_{z}$ chaotic orbits is that they fill the region they traverse in $(l_{\theta,\rm p},\theta_{\rm p})$ uniformly (the same is true at apoapse, though that is less crucial for the following analyses). This behaviour is shown in Figure 6 where we integrate a range of trajectories over 100,000 radial periods. From this we count the number of times any given periapse has a certain $\theta_{\rm p}$ and $l_{\theta,\rm p}$.
From left to right, the panels show trajectories of increasing $l_{z}$, in the range $0.01\leq l_{z}/i_{c}\leq 0.1$. These low-$l_{z}$ trajectories occupy a complex region, that is bounded from above in $l_{\theta,\rm p}$ and also avoids islands of regular orbits. Some of these excluded regions may be fractal in nature (corresponding to harmonics of the dominant local regular orbit) and thus smaller than what is resolved by this plot, while other parts of the chaotic region may be narrow enough to not span an entire cell - all of which can lead to a decreased number density. Nevertheless, by visually examining the whole parameter space of low-$l_{z}$ trajectories, we see that the occupation fraction within the chaotic sea generally appears uniform. To make this analysis quantitative, we take the counts from the plots and show their Cumulative Distribution Function (CDF). We compare each trajectory’s CDF to that of a Poisson distribution, which we would expect if all cells contained the same number of counts $N_{\rm c}$ and varied only by shot noise $\sqrt{N_{\rm c}}$. We see that the Poisson distribution agrees very well with the observed counts at smaller $l_{z}$. Generally speaking, we find excellent agreement with the Poisson distribution (and thus uniform occupancy of the chaotic region) for $l_{z}/l_{\rm c}\lesssim 0.1$. Cells spanning an edge of the chaotic region may have a slightly diminished count, leading to an excess of low counts - as we see in some panels, where the red line exceeding the Poisson distribution at low counts. However other boundaries display a ’sticky’ behaviour (for example in the top middle panel) with trajectories tracing the edge of a regular region for many orbits before diverging.
@@
One interesting feature, most apparent in the second cell of the first row ($R_{\rm c}=1$ pc, $l_{z}=0.03\ l_{\rm c}$), is that some embedded regular orbits seem to act as attractors - with the chaotic trajectory visiting their periphery much more frequently than elsewhere. In the second panel of the second row ($R_{\rm c}=3$ pc, $l_{z}=0.03\ l_{\rm c}$) we observe another behaviour: a minor but significant drop in density in one region (associated with low $l_{\theta,\rm p}$ and intermediate $\theta_{\rm p}$). Within this region there are also a series of high density points, again likely corresponding to the region around a (higher harmonic) regular orbit.
@@ 4.4 Diving orbits
Figure 7: Top panel: the fraction of trajectories which dive to arbitrarily small periapsides ($F_{\rm dive}$). Middle panel: the fraction of chaotic trajectories that do not dive ($1-\frac{F_{\rm dive}}{F_{\rm chaos}}$). Bottom panel: the normalising fraction of bound trajectories ($F_{\rm phys}$). The top two panels are found by simulating 100,000 trajectories at random $R_{\rm c}$ and $l_{z}/l_{\rm c}$ values (uniformly sampled in the space shown), initialized with uniform random $l_{\theta,\rm p}/l_{\rm c}$ and $\theta_{\rm p}$, and integrated for approximately 400 orbits each. The bottom panel is computed from equation 27. We see that an diving orbits are almost always a majority of chaotic orbits.
@@
Not every chaotic trajectory comes as close as possible to the MPP, as some chaotic trajectories occupy the definitely rotating region of the SoS. We use a simple criterion to separate these less relevant chaotic trajectories from the diving trajectories of interest, namely that $\min(|l_{\theta,\rm p}/l_{\rm c}|)<N_{\rm p}^{-\frac{1}{2}}$ where $N_{\rm p}$ is the number of periapse passages. This criterion is motivated by the assumption of uniform occupation of the chaotic region, meaning that if the region extends to $l_{\theta,\rm p}=0$ the minimum observed value will be small, but limited by Poisson noise (hence the comparison to $N_{\rm p}^{-\frac{1}{2}}$). This ignores the detailed shape and size of the full chaotic region, but we have found that this simple criterion cleanly separates diving orbits from confined chaotic orbits. One caveat is that a reliable measure requires $N_{\rm p}\gg 1$; we will mostly use 400 or more periapse passages for our analysis.
@@
Near the poles of a flattened potential, some values of $|l_{\theta}|<l_{\rm c}$ may result in unbound trajectories [^9] We remind the reader that $l_{\rm c}$ is the maximum angular momentum for the whole system but that it corresponds to orbits in the equatorial plane, where the potential is itself maximzed., i.e. those with initial $l_{\theta}>l_{\theta,\rm max}(\theta)$ (see equation 18). We can discard these unphysical orbits in two different ways. The efficient but more complex way is to choose random $\theta$ values, weighted by $l_{\theta,\rm max}(\theta)/l_{\rm c}$, and then to choose uniform $|l_{\theta}|<l_{\theta,\rm max}(\theta)$. The simpler method is to choose uniform random $\theta$ and $|l_{\theta}|<l_{\rm c}$ and then reject initial conditions with $|l_{\theta}|>l_{\theta,\rm max}(\theta)$. We adopt the latter approach, and renormalise our fractions by those of physical orbits,
@@
We see that the fraction of non-diving chaotic orbits is small, at most around a third, but only reaching these values at large $R_{\rm c}$ ($>100$ pc). Thus we can conclude that the vast majority of chaotic orbits lead to (or, as discussed in section 4.2, stem from) arbitrarily close periapse passages.
@@
4.5 Comparing to Magorrian and Tremaine 1999
Our analysis is partly informed by the calculations presented in Magorrian & Tremaine (1999), in the sections where they assess the rate of disruptions caused by precession of orbits in axisymmetric potentials. They too utilise SoSs to estimate the region of parameter space where chaotic orbits dominate. Our work differs from theirs in one minor facet and one major respect, and thus our eventual results also differ.
Magorrian & Tremaine (1999) produce and draw inferences from a SoS, though only at apoapse. Unlike the SoSs used in our work, which we calculate from direct orbit integration, theirs comes from a discrete, orbit-averaged (Chirikov-like) map, which as in [^10] Originally this method was developed for slightly triaxial systems to model precession in the azimuthal angle, but Magorrian & Tremaine (1999) apply it to the polar angle in the axisymmetric case. Touma & Tremaine (1997) is used to calculate the successive $\theta_{\rm a},\ l_{\theta,\rm a}$ values. We have compared the two approaches (direct integration versus the orbit-averaged discrete map) and found that while the Magorrian & Tremaine (1999) approach works well in the limit of perturbative axisymmetry, it breaks down for non-perturbative deviations from sphericity. The orbit-averaged method assumes that the variation in state between successive apoapses are small, implicitly relying on the assumption of mild axisymmetry, which is not satisfied in our GC model. We futher note that the Magorrian & Tremaine (1999) SoS analysis is not directly used in their later calculations of rates, but only to more fully inform the discussion, whearas the fraction and properties of chaotic regions in the SoS will directly inform our subsequent analysis.
The larger difference between these two works is that Magorrian & Tremaine (1999) calculate rates based on the fraction of total time which trajectories spend with close to zero angular momentum, wheareas our results will only use the fraction of periapse passages with $l\rightarrow 0$. These fractions are quite different, and the latter is the one directly relevant to disruptions (zero angular momenta at large radii correspond to a radial trajectory, but not necessarily a disruption). Propagating these different assumptions through to timescales and rates of disruptions (as we will do across the next two sections) leads to quite different inferences about the rates and properties of disruptions.
In most of the galaxies considered by Magorrian & Tremaine (1999), they find a moderate (order of magnitude or less) increase to the overall rates of disruptions when including the collisionless channel. We will compare to this finding in Section 7.
@@ 4.6 Varying the NSC model
A point mass alone produces closed Keplerian orbits; adding a sub-dominant spherical mass distribution (or including relativistic effects) will generate precession, but will not produce chaotic behaviour. It is the non-sphericity of our simulated GC that gives rise to chaos, and thus in this section we primarily investigate the effect of the degree of flattening of the NSC. This is encapsulated in the parameter $q_{\rm NSC}$. We use a slightly simplified version of our GC potential for these tests, by defining the full potential rather than an initial guess (which is then iterated upon to fit a fixed NSD DF) as in the main work.
@@
where $M_{\rm BH}=4\cdot 10^{6}\ M_{\odot}$ is the BH mass and $a_{\rm BH}=10^{-3}$ pc is the scale radius (i.e. the BH potential is slightly softened for numerical stability, but only on very small scales below the range our orbits probe [^11] We note that $a_{\rm BH}$ is often larger than the tidal separation radius for the Hills mechanism, and is always much larger than the disruption radius for TDEs, but remind the concerned reader that our orbital integrations are performed primarily to characterize the topology and density of low-$l_{z}$ chaotic orbits, not to simulate disruptive processes directly.). The NSC as in the fiducial model of S22: as a density profile (based on a fit to data from Chatzopoulos et al. 2015) which is then converted to a potential by AGAMA. More specifically, it is defined via the axisymmetric generalisation of a Dehnen profile (Dehnen, 1993; Merritt & Fridman, 1996)
@@
We repeat an analysis similar to the one used in Figure 8. We limit our interest to orbits with $l_{z}=0$. For each $R_{\rm c}$, we initialise 100 orbits with random $l_{\theta}$ and $\theta$ at periapse, and follow them for $\approx$400 radial periods. We check whether orbits are physical (as some regions of $l_{\theta},\theta$ space are not accessible in flattened potentials) and chaotic.
@@
Figure 9: Top panel: The fraction of chaotic orbits (solid line) and physical orbits (dashed line) for a given NSC axis ratio ($q_{\rm NSC}$) as a function of $R_{\rm c}$. Bottom panel: The same but in a GC model with only BH and NSC (no NSD). We see that realistically flattened NSCs produce order-unity chaotic fractions, and even nearly spherical NSCs retain some chaotic orbits. The NSD increases the chaotic fraction at distances $R_{\rm c}\gtrsim 5$ pc.
@@ 5 Probabilities, properties and timescale of diving orbit disruptions
The uniform occupation of trajectories across $l_{\theta,\rm p},\theta_{\rm p}$ space is a strong assumption, which may not in general be true (see for example Thomas et al. 2004), even as it remains true for any single chaotic trajectory. Brief numerical experiments show that the DFs of our GC model prefer low $|l_{\theta,\rm p}|$ and $\theta_{\rm p}$ close to $\frac{\pi}{2}$. This would boost the effective $F_{\rm dive}$ as this is the region where diving orbits are most frequently found. This makes our simplifying assumption a conservative one.
@@
which also runs from 0 to 1 . They are both uniformly sampled by periapse passages. In many cases the transformed coordinate
@@
which is 0 at the poles and 1 at the equator, will be even more expedient to use. In this notation we can write
@@ 5.1.2 Normalizing $p(\beta)$
Here we have assumed a single definite $l_{\rm t}$ that can separate between which encounters lead to disruptions. These simplifying assumption can be complicated for Hills mechanism separations, where disruptions can occur for initially moderate $\beta$ (e.g. $\beta\gtrsim 0.5$ as shown in Sari et al. 2010), and a minority of systems may survive encounters with $\beta\gg 1$ (see again Sari et al. 2010 and Sersante et al. 2025, in prep.).
A fuller description of the loss cone dynamics could instead take a probabilistic approach, but that is beyond the scope of this work.
@@ 5.3 Probability and timescales
$p(D)=p(D|\rho,\gamma_{z},l_{\rm t})=\begin{cases}\lambda_{\rm t}\frac{1-|\gamma_{z}|}{F_{\rm dive}(\rho)}&|\gamma_{z}|\leq 1\\ 0&\text{otherwise}\end{cases}$
@@ 6 Rates of disruption for a given system
In this section, we use our earlier dynamical results to estimate the rate at which systems (e.g. stars or binaries) are delivered to small pericenters where they may be disrupted. In a spherically symmetric potential, the small 3-dimensional volume of $l$-space (with $l\leq l_{\rm t}$; Frank & Rees 1976) vulnerable to disruption is called the “loss cone,” which may be either empty or full depending on the ratio of a collisional relaxation time to the orbital time. In an axisymmetric potential such as our model of the Milky Way, the larger 1-dimensional volume of $\mathbf{l}$-space vulnerable to disruption is called the “loss wedge” (defined by $l_{z}\leq l_{\rm t}$; Magorrian & Tremaine 1999) which in turn may be either empty or full (as detailed in section 6.2. Collisionless motion of trajectories through the loss wedge will often be efficient at delivering objects to small pericenters. While this motion is sometimes due to libration along regular orbits (Vasiliev & Merritt, 2013), our earlier results (section 4) show that for a realistic Milky Way model, chaotic orbits overwhelmingly dominate collisionless disruption rates from the loss wedge.
Our calculations so far have taken $l_{\rm t}$ as a given. To fully calculate rates we need to know not just what happens to a system with fixed $l_{\rm t}$, but also how numerous such systems are compared to other systems with different $l_{\rm t}$. We save this multi-species calculation for an upcoming paper (Penoyre et al. 2025, in prep.) where we will consider both primordial populations of systems and the condition that they must survive long enough to disrupt. For the rest of this paper we present the calculation of rates assuming a single value of $l_{\rm t}$, thus giving a qualitative understanding of the total rate of disruptions we might expect, as well as quantitative predictions for the relative rates of disruptions from various mechanisms. In particular, alongside the collisionless rate calculated for diving orbits we will also compute (a simple version of) the relative collisional rate expected from scattering events, and directly compare these.
$\Gamma_{\rm i}$ is the rate of disruption from a given mechanism (where $\rm i$ is either $\rm d$ for diving orbits or $\rm s$ for scattering). We can write a general simplified form of the rate depending on some set of properties, which we can express as a vector $\vec{q}_{\rm i}$ of length $N$, giving
@@
where $d\mathcal{N}=n(\vec{q}_{\rm i})d\vec{q}_{\rm i}$ is the differential number of systems with properties in the range $\vec{q}_{\rm i}$ to $\vec{q}_{\rm i}+d\vec{q}_{\rm i}$, and $P_{\rm i}(D|\vec{q}_{\rm i})$ and $t_{\rm i}(\vec{q}_{\rm i})$ are the probability and the expected timescale respectively that such a system disrupts by mechanism $\rm i$. To find the total disruption rate, we would need to integrate over each individual property in $\vec{q}_{\rm i}$ in turn. For diving orbits we have shown that the relevant properties in $\vec{q}_{\rm d}$ are $(\epsilon,\ l_{z},\ l_{\rm t})$ (or transformations thereof). For scatterings (assuming an isotropic velocity distribution) the equivalent properties are $\vec{q}_{\rm s}=(\epsilon,\ l,\ l_{\rm t})$, dependent on the total angular momentum not just the z-component.
@@
where $l_{\rm d}=l_{z}$ and $l_{\rm s}=l$. In this form, $\bar{\Gamma}$ is equivalent to the rate of disruptions that would be observed if every system in the model had this single identical value of $l_{\rm t}$.
There is in general no direct relationship between $\bar{\Gamma}_{\rm i}$ and $\Gamma_{\rm i}$ (or any of their derivatives). We might hope that $\bar{\Gamma}_{\rm i}(l_{\rm t})=a(l_{\rm t})\frac{d\Gamma_{\rm i}}{dl_{\rm t}}$ (i.e. that we could find the true rate, $\Gamma_{\rm i}$, by integrating some function of $\bar{\Gamma}_{\rm i}$ over $l_{\rm t}$) but there are two strong conditions that determine if this is possible.
@@
  • The integral of equation 60 spans the range $|l_{\rm i}|\leq l_{\rm t}$, thus for these integrals to not introduce further irreconcilable factors of $l_{\rm t}$ they must depend negligibly on their upper limit. In other words, almost all disruptions have $|l_{\rm i}|\ll l_{\rm t}$

### v3

@@ 4.2.1 Lyapunov exponents
$\lambda$ should only be calculated for the section of a trajectory for which the displacement is growing exponentially. It is common to use the convention of $\lambda=0$ for trajectories which show no such exponential growth and thus are assumed to be regular. Following the scheme outlined in Vasiliev (2013), we can efficiently estimate $\lambda$ in units of the circular period $T_{\rm c}$, such that its value is normally between 0 and a maximum close to 1 (corresponding to an trajectory that diverges over a single orbit).
A completely general and robust quantification of $\lambda$ is a difficult task and we note two edge cases that infrequently occur. [^9] These issues have been fixed in the most recent update to AGAMA, however our results are still based on analysis using an older version of the code where they are still present. The first comes from demanding too high an integration accuracy from the integrator, which can result in ’ghost’ timesteps (with ${\rm d}t_{i}=t_{i+1}-t_{i}=0$) causing $\lambda$ to be undefined. The second comes from lower integration accuracy, leading to larger timesteps which are still well optimized for resolving the orbital path but which may now be insufficiently spaced for calculating $\lambda$, and thus give erroneous $\lambda>0$ for regular orbits (as identified, for example from the SoS).
@@ 4.3 Uniform filling of the chaotic region
[…]
From left to right, the panels show trajectories of increasing $l_{z}$, in the range $0.01\leq l_{z}/i_{c}\leq 0.1$. These low-$l_{z}$ trajectories occupy a complex region, that is bounded from above in $l_{\theta,\rm p}$ and also avoids islands of regular orbits. Some of these excluded regions may be fractal in nature (corresponding to harmonics of the dominant local regular orbit) and thus smaller than what is resolved by this plot, while other parts of the chaotic region may be narrow enough to not span an entire cell - all of which can lead to a decreased number density. Nevertheless, by visually examining the whole parameter space of low-$l_{z}$ trajectories, we see that the occupation fraction within the chaotic sea is close to uniform. To make this analysis quantitative, we take the counts from the plots and show their Cumulative Distribution Function (CDF). We compare each trajectory’s CDF to that of a Poisson distribution, which we would expect if all cells contained the same number of counts $N_{\rm c}$ and varied only by shot noise $\sqrt{N_{\rm c}}$. We see that the Poisson distribution agrees very well with the observed counts at smaller $l_{z}$. Generally speaking, we find excellent agreement with the Poisson distribution (and thus uniform occupancy of the chaotic region) for $l_{z}/l_{\rm c}\lesssim 0.1$. Cells spanning an edge of the chaotic region may have a slightly diminished count, leading to an excess of low counts - as we see in some panels, where the red line exceeding the Poisson distribution at low counts.
Other boundaries display a ‘sticky’ behaviour (for example in the top middle panel) with trajectories oversampling the edge of a regular region. This may be due to numerical integration error allowing the trajectory to drift into a regular region for a while, completing a number of orbits before drifting back into the chaotic region.
@@ 4.4 Diving orbits
Figure 7: Top panel: the fraction ($F_{\rm dive}$) of trajectories which dive to arbitrarily small $l_{\theta,p}$ and thus reach the minimum pericenter permitted by $l_{z}$ conservation. Middle panel: the fraction of chaotic trajectories that do not dive ($1-\frac{F_{\rm dive}}{F_{\rm chaos}}$). Bottom panel: the normalising fraction of bound trajectories ($F_{\rm phys}$). The top two panels are found by simulating 100,000 trajectories at random $R_{\rm c}$ and $l_{z}/l_{\rm c}$ values (uniformly sampled in the space shown), initialized with uniform random $l_{\theta,\rm p}/l_{\rm c}$ and $\theta_{\rm p}$, and integrated for approximately 400 orbits each. We see that diving orbits are almost always a majority of chaotic orbits.
@@
Not every chaotic trajectory comes as close as possible to the MPP, as some chaotic trajectories occupy the definitely rotating region of the SoS. We use a simple criterion to separate these less relevant chaotic trajectories from the diving trajectories of interest, namely that $\min(|l_{\theta,\rm p}/l_{\rm c}|)<N_{\rm p}^{-\frac{1}{2}}$ where $N_{\rm p}$ is the number of periapse passages. This criterion is motivated by the assumption of ergodicity (uniform occupation of the chaotic region) meaning that if the region extends to $l_{\theta,\rm p}=0$ the minimum observed value will be small, but limited by Poisson noise (hence the comparison to $N_{\rm p}^{-\frac{1}{2}}$). This ignores the detailed shape and size of the full chaotic region, but we have found that this simple criterion cleanly separates diving orbits from confined chaotic orbits. One caveat is that a reliable measure requires $N_{\rm p}\gg 1$; we will mostly use 400 or more periapse passages for our analysis.
@@
Near the poles of a flattened potential, some values of $|l_{\theta}|<l_{\rm c}$ may result in unbound trajectories, [^10] We remind the reader that $l_{\rm c}$ is the maximum angular momentum for the whole system but that it corresponds to orbits in the equatorial plane, where the potential is itself maximzed. i.e. those with initial $l_{\theta}>l_{\theta,\rm max}(\theta)$ (see equation 18). We can discard these unphysical orbits in two different ways. The efficient but more complex way is to choose random $\theta$ values, weighted by $l_{\theta,\rm max}(\theta)/l_{\rm c}$, and then to choose uniform $|l_{\theta}|<l_{\theta,\rm max}(\theta)$. The simpler method is to choose uniform random $\theta$ and $|l_{\theta}|<l_{\rm c}$ and then reject initial conditions with $|l_{\theta}|>l_{\theta,\rm max}(\theta)$. We adopt the latter approach, and renormalise our fractions by those of physical orbits,
@@
We see that the fraction of non-diving chaotic orbits (as shown in the lower panel) is small, at most around a third, but only reaching these values at large $R_{\rm c}$ ($>100$ pc). Thus we can conclude that the vast majority of chaotic orbits lead to (or, as discussed in section 4.2, stem from) arbitrarily close periapse passages.
@@ 4.5 Varying the NSC model
A point mass alone produces closed Keplerian orbits. Adding a sub-dominant spherical mass distribution (or including relativistic effects) will generate precession, but will not produce chaotic behaviour. It is the non-sphericity of our simulated GC that gives rise to chaos, and thus in this section we primarily investigate the effect of the degree of flattening of the NSC. This is encapsulated in the parameter $q_{\rm NSC}$.
@@
where $M_{\rm BH}=4\cdot 10^{6}\ M_{\odot}$ is the BH mass and $a_{\rm BH}=10^{-3}$ pc is the scale radius (i.e. the BH potential is slightly softened for numerical stability, but only on small scales below the range most of our orbits probe [^11] We note that $a_{\rm BH}$ is often larger than the tidal separation radius for the Hills mechanism, and is always much larger than the disruption radius for TDEs, but remind the concerned reader that our orbital integrations are performed primarily to characterize the topology and density of low-$l_{z}$ chaotic orbits, not to simulate disruptive processes directly.). The NSC as in the fiducial model of S22: as a density profile (based on a fit to data from Chatzopoulos et al. 2015) which is then converted to a potential by AGAMA. More specifically, it is defined via the axisymmetric generalisation of a Dehnen profile (Dehnen, 1993; Merritt & Fridman, 1996)
@@
We repeat an analysis similar to the one used in Figure 8. We limit our interest to orbits with $l_{z}=0$. For each $R_{\rm c}$, we initialise 100 orbits with random $l_{\theta}$ and $\theta$ at periapse, and follow them for $\approx$400 radial periods. We check whether orbits are physical (as some regions of $l_{\theta},\theta$ space are not accessible in flattened potentials) and chaotic, equivalent to $F_{\rm phys}$ and $F_{\rm chaos}$ from the previous section.
@@
Figure 9: Top panel: The fraction of chaotic orbits ($F_{\rm chaos}$, solid line) and physical orbits ($F_{\rm phys}$, dashed line) for a given NSC axis ratio ($q_{\rm NSC}$) as a function of $R_{\rm c}$. Bottom panel: The same but in a GC model with only BH and NSC (no NSD). We see that realistically flattened NSCs produce order-unity chaotic fractions, and even nearly spherical NSCs retain some chaotic orbits. The NSD increases the chaotic fraction at distances $R_{\rm c}\gtrsim 5$ pc.
@@ 5 Probabilities, properties and timescale of diving orbit disruptions
The strong assumption of ergodicity, uniform occupation of trajectories across $l_{\theta,\rm p},\theta_{\rm p}$ space, may not in general be true (see for example Thomas et al. 2004), even as it remains true for any single chaotic trajectory. Brief numerical experiments show that the DFs of our GC model prefer low $|l_{\theta,\rm p}|$ and $\theta_{\rm p}$ close to $\frac{\pi}{2}$. This would boost the effective $F_{\rm dive}$ as this is the region where diving orbits are most frequently found. This makes our simplifying assumption a conservative one.
@@
which also runs from 0 to 1 . They are both uniformly sampled by periapse passages. In many cases it is expedient to use the transformed coordinate
@@
which is 0 at the poles and 1 at the equator. In this notation we can write
@@ 5.1.2 Normalizing $p(\beta)$
Here we have assumed a single definite $l_{\rm t}$ that can separate between which encounters lead to disruptions. These simplifying assumption can be complicated for Hills mechanism separations, where disruptions can occur for initially moderate $\beta$ (e.g. $\beta\gtrsim 0.5$), and a minority of systems may survive encounters with $\beta\gg 1$ (see Sari et al. 2010 and Sersante et al. 2025).
A fuller description of the loss-cone dynamics could instead take a probabilistic approach, but that is beyond the scope of this work.
@@ 5.3 Probability and timescales
$p(D)=p(D|\rho,\gamma_{z},l_{\rm t})=\begin{cases}\lambda_{\rm t}\frac{1-|\gamma_{z}|}{F_{\rm dive}(\rho)}&|\gamma_{z}|\leq 1\\ 0&\text{otherwise}\end{cases}$ (57)
@@ 6 Rates of disruption for a given system
In this section, we use our earlier dynamical results to estimate the rate at which systems (e.g. stars or binaries) are delivered to small pericenters where they may be disrupted. In a spherically symmetric potential, the small 3-dimensional volume of $\mathbf{l}$-space (with $l\leq l_{\rm t}$; Frank & Rees 1976) vulnerable to disruption is called the “loss cone,” which may be either empty or full depending on the ratio of a collisional relaxation time to the orbital time. In an axisymmetric potential such as our model of the Milky Way, the larger 1-dimensional volume of $\mathbf{l}$-space vulnerable to disruption is called the “loss wedge” (defined by $|l_{z}|\leq l_{\rm t}$; Magorrian & Tremaine 1999) which in turn may be either empty or full (as detailed in section 6.2). Diving orbits, collisionless trajectories that pass through the loss wedge, can be efficient at delivering objects to small pericenters. While this motion is sometimes due to libration along regular orbits (Vasiliev & Merritt, 2013), our earlier results (section 4) show that for a realistic Milky Way model, chaotic orbits overwhelmingly dominate collisionless disruption rates from the loss wedge.
Our calculations so far have taken $l_{\rm t}$ as a given. To fully calculate rates we need to know not just what happens to a system with fixed $l_{\rm t}$, but also how numerous such systems are compared to other systems with different $l_{\rm t}$. We save this multi-species calculation for an upcoming paper (Penoyre et al. 2025, in prep.) where we will consider both primordial populations of systems and the condition that they must survive long enough to disrupt. For the rest of this paper we present the calculation of rates assuming a single value of $l_{\rm t}$, thus giving a qualitative understanding of the total rate of disruptions we might expect, as well as quantitative predictions for the relative rates of disruptions from various mechanisms. In particular we will calculate and compare a simplified collisionless rate, from diving orbits, and collisional rate, from scattering events.
$\Gamma_{\rm i}$ is the rate of disruption from a given mechanism (where $\rm i$ is either $\rm d$ for diving orbits or $\rm s$ for scattering). Any rate depends on some set of properties, which we can express as a vector $\vec{q}_{\rm i}$ of length $N$, and can be written as
@@
where $d\mathcal{N}=n(\vec{q}_{\rm i})d\vec{q}_{\rm i}$ is the differential number of systems with properties in the range $\vec{q}_{\rm i}$ to $\vec{q}_{\rm i}+d\vec{q}_{\rm i}$, and $P_{\rm i}(D|\vec{q}_{\rm i})$ and $t_{\rm i}(\vec{q}_{\rm i})$ are the probability and the expected timescale respectively that such a system disrupts by mechanism $\rm i$. To find the total disruption rate, we would need to integrate over each individual property in $\vec{q}_{\rm i}$ in turn. For diving orbits we have shown that the relevant properties in $\vec{q}_{\rm d}$ are $(\epsilon,\ l_{z},\ l_{\rm t})$ (or transformations thereof). For scatterings the equivalent properties are $\vec{q}_{\rm s}=(\epsilon,\ l,\ l_{\rm t})$, dependent on the total angular momentum not just the z-component.
@@
In this form, $\bar{\Gamma}$ is equivalent to the rate of disruptions that would be observed if every system in the model had this single identical value of $l_{\rm t}$.
There is in general no direct relationship between $\bar{\Gamma}_{\rm i}$ and $\Gamma_{\rm i}$ (or any of their derivatives). We might hope that $\bar{\Gamma}_{\rm i}(l_{\rm t})=a(l_{\rm t})\frac{d\Gamma_{\rm i}}{dl_{\rm t}}$ (i.e. that we could find the true rate, $\Gamma_{\rm i}$, by integrating some function of $\bar{\Gamma}_{\rm i}$ over $l_{\rm t}$) but there are two strong conditions that must be met for this to be possible.
@@
  • The integral over equation 61 spans the range $|l_{\rm i}|\leq l_{\rm t}$, thus for these integrals to not introduce further irreconcilable factors of $l_{\rm t}$ they must depend negligibly on their upper limit. In other words, almost all disruptions have $|l_{\rm i}|\ll l_{\rm t}$